Let us consider two irreducible fractions. The denominator of the first one is equal to 8200,and the denominator of the second to 4300. What is the smallest possible denominator of a fraction equal to the sum of these fractions, after the fraction is reduced? (For example, (2/3) + (8/15) = (18/15) = (6/5), and the denominator after the reduction is equal to 5.)
1
Expert's answer
2020-12-02T12:02:18-0500

We have

"\frac{x}{8200}+\frac{y}{4300}=\frac{43x+82y}{352600} \\ \text{But the gcd(43x+82y,352600)} = \frac{(43x+82y)352600}{lcm(43x+82y,352600)}, \text{ then for every } x,y\in \mathbb{Z} \text{ gcd(43x+82y,352600) divides (43x+82y) and 352600} \\ \therefore \boxed{\text{ the lowest denominator is } \frac{352600}{gcd(43x+82y,352600)}=\frac{lcm(43x+82y,352600)}{43x+82y} } \\ \text{For example if x=y=1 then we will have the lowest denominator to be } \frac{lcm(43+82,352600)}{43+82}= \frac{lcm(125,352600)}{125}=\frac{1763000}{125}=14104"